if AX =45 find the value of each of the following.
2y+1 | y-1
.__________._____________.
A Q X
|---------45-------------|

Answer by cleomenius(959)

AQ = 2y + 1
Qx = y - 1
2y + 1 + y - 1 = 45
3y = 45
y = 15
AQ = 31
QX = 14
31 + 14 = 45.
This does check.
Cleomenius.